OpenStudy (anonymous):

Solve this system of linear equations. Separate the x- and y-values with a comma. -6x+2y=4 -5x+5y=0

OpenStudy (anonymous):

1,1

OpenStudy (anonymous):

for the second equation, we can add -5x to both sides, resulting in the equality 5y=5x dividing by 5 on both sides, we get the equality y=x now, we plug this into the first equation, telling us that -6x+2x=4 putting the variables together -4x=4 -x=1 x=-1 and since x=y y=-1 so the answer is -1, -1